### v1

# Extracting the Luttinger parameter from a single wave function

Abstract
The low-energy physics of Tomonaga-Luttinger liquids (TLLs) is controlled by the Luttinger parameter. We demonstrate that this parameter can be extracted from a single wave function for one-component TLLs with periodic boundary condition. This method relies on the fact that TLLs are described by conformal field theory in which crosscap states can be constructed. The overlaps between the crosscap states and the ground state as well as some excited states are proved to be universal numbers that directly reveal the Luttinger parameter. In microscopic lattice models, crosscap states are formed by putting each pair of antipodal sites into a maximally entangled state. Analytical and numerical calculations are performed in a few representative models to substantiate the conformal field theory prediction. The extracted Luttinger parameters are generally quite accurate in finite-size systems with moderate lengths, so there is no need to perform data fitting and/or finite-size scaling.
Introduction — The theory of Tomonaga-Luttinger liquid (TLL) is a great triumph of strongly correlated physics [1]. Based on previous work of Tomonaga [2], Luttinger introduced a model to study interacting fermions in one dimension (1D) [3]. It clearly exemplifies the peculiarity of reduced dimensionality because infinitesimal interaction in a 1D system drives it away from the Fermi liquid. In subsequent works [4, 5, 6, 7], the bosonization framework was established to provide a unified treatment for many strongly correlated problems in 1D. Spin chains are usually studied using the Jordan-Wigner transformation that converts spins to fermions. Gapless phases of bosons or fermions can be understood using free boson fields that correspond to density fluctuations. Instabilities of the TLL lead to symmetry-breaking ordered phases. It is not easy to realize 1D systems, but experimental investigations have been carried out in solid state and cold atom platforms [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Another context in which TLL thrives is the edge of certain two-dimensional topological states [18, 19, 20, 21].
The low-energy physics of one-component TLLs is captured by the Hamiltonian
$H=\frac{v}{8\pi}\int^{L}_{0}\mathrm{d}x\,\left[(\partial_{x}\hat{\varphi})^{2}+(\partial_{x}\hat{\theta})^{2}\right].$ (1)
In the conformal field theory (CFT) language, this is a system of compactified bosons. Here $L$ is the length of the system, $v$ is the characteristic velocity, $\hat{\varphi}$ is a compactified boson field with radius $R$ (i.e., $\hat{\varphi}\sim\hat{\varphi}+2\pi R$), and $\hat{\theta}$ is the conjugate boson field of $\hat{\varphi}$ with radius $R^{\prime}=2/R$. The Luttinger parameter $K$ can be expressed using the radius as $K=R^{2}/4$. Its value is $1$ for free fermions and deviates from $1$ when interaction is turned on. For a given TLL, many properties are determined by $K$ so it is routinely computed in numerical studies. If the Hamiltonian has a microscopic U(1) symmetry (e.g., particle number conservation), $K$ can be obtained by comparing the ground-state energy of multiple systems with different U(1) charges [22]. This is feasible because $K$ can be interpreted as the stiffness of a superfluid. The powerlaw decay of certain correlation functions can be used to deduce $K$ when their bosonized expressions are known [1]. It has been found that $K$ also appears in the U(1) charge fluctuations of a subsystem [23, 24], the U(1) symmetry-resolved entanglement spectrum [25, 26], or the entanglement entropy associated with particle partition [27, 28].
In this Letter, we propose that the Luttinger parameter in microscopic models can be extracted from the eigenstate crosscap overlap (ESCO). It has three favorable features: i) a single wave function instead of multiple states is used; ii) data fitting or finite-size scaling is not needed; iii) microscopic U(1) symmetry is not required. To the best of our knowledge, no other method enjoys all three advantages simultaneously. Our method is motivated by classical results and recent progress in CFT and integrable models. When a CFT is defined on the Klein bottle, important information can be inferred from the thermal entropy [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. The Klein bottle can be viewed as a special cylinder with crosscaps at its ends. A simple circle would turn into a crosscap if all pairs of antipodal points are identified. It is possible to define a variety of crosscap states in the framework of boundary CFT [42]. For a rational CFT, the crosscap states have universal overlaps (“crosscap coefficients”) with its ground state [43, 44, 45]. In some integrable models, the crosscap states have been identified as “integrable boundary states” [46, 47, 48, 49, 50]. The overlaps between all eigenstates of these integrable models and the crosscap states can be expressed as determinants, but their physical significance is not transparent. We aim to establish a general framework for studying crosscap states in CFT (without and with perturbations) such that physical properties can be deduced from their overlaps with eigenstates.
Figure 1: (a) Schematics of the modified Villain model and its crosscap states $|\mathcal{C}_{\pm}\rangle$. (b) Graphic representations of an MPS with eight sites and its crosscap overlap.
General construction of crosscap states — While crosscap is simple to visualize in the continuum, crosscap states in lattice models have only been constructed in a somewhat intuitive manner. The main obstacle is how to make a connection between the degrees of freedom in the continuum and lattice. To this end, we find it particularly instructive to use the modified Villain model [51]
$H_{\mathrm{mV}}=\sum_{j=1}^{N}\left[(2\pi R\,\hat{p}_{j})^{2}+{R^{\prime}}^{2}(\hat{X}_{j+1}-\hat{X}_{j}+2\pi\hat{n}_{j,j+1})^{2}\right]\,,$ (2)
as illustrated in Fig. 1(a). It provides a specific lattice realization of the compactified boson Hamiltonian in Eq. (1). There are $N$ sites in the chain (labeled by $j$), the lattice spacing is $a$, and periodic boundary condition is imposed. $\hat{X}_{j}$ ($\hat{E}_{j,j+1}$) is the site (link) variable at site $j$ [link $(j,j+1)$] and $\hat{p}_{j}$ ($\hat{n}_{j,j+1}$) is its canonical momentum. The standard commutation relations $[\hat{X}_{j},\hat{p}_{l}]=[\hat{E}_{j,j+1},\hat{n}_{l,l+1}]=i\delta_{jl}$ are satisfied. The local Hilbert space at site $j$ [link $(j,j+1)$], denoted as $\mathcal{H}_{j}$ ($\mathcal{H}_{j,j+1}$), is spanned by the eigenbasis of $\hat{X}_{j}$ ($\hat{E}_{j,j+1}$) and expressed as $|X_{j}\rangle$ with $X_{j}\in\mathbb{R}$ ($|E_{j,j+1}\rangle$ with $E_{j,j+1}\in[0,2\pi)$). The full Hilbert space is thus the tensor product of local ones on all sites and links, $\mathcal{H}=\prod^{N}_{j=1}\mathcal{H}_{j}\otimes\mathcal{H}_{j,j+1}$. However, local operators $\hat{G}_{j}=e^{i(\hat{E}_{j,j+1}-\hat{E}_{j-1,j}-2\pi\hat{p}_{j})}$ commute with both terms in the Hamiltonian (2), $[\hat{G}_{j},\hat{p}_{j^{\prime}}]=[\hat{G}_{j},\hat{X}_{j^{\prime}+1}-\hat{X}_{j^{\prime}}+2\pi\hat{n}_{j^{\prime},j^{\prime}+1}]=0\;\forall j,j^{\prime}$, and hence also commute with the Hamiltonian, $[H_{\mathrm{mV}},\hat{G}_{j}]=0\;\forall j$. Thus, we may impose the Gauss law constraint $\hat{G}_{j}=1$ such that the model becomes a lattice gauge theory. The physical Hilbert space is the subspace $\mathcal{H}_{\mathrm{phy}}$ of the full space $\mathcal{H}$ that satisfies the Gauss law. The projection operator onto the subspace $\mathcal{H}_{\mathrm{phy}}$ is denoted as $\mathcal{P}_{G}$.
To transform the model from lattice to continuum, we introduce new variables $\hat{\theta}_{j}=R^{\prime}(\hat{X}_{j}+2\pi\sum_{k=1}^{j}\hat{n}_{k-1,k})$ and $\hat{\varphi}_{j,j+1}=R\hat{E}_{j,j+1}$ [51]. It is obvious that $[e^{i\hat{\theta}_{j}/R^{\prime}},\hat{G}_{j^{\prime}}]=[e^{i\hat{\varphi}_{j,j+1}/R},\hat{G}_{j^{\prime}}]=0\;\forall j,j^{\prime}$, so $e^{i\hat{\theta}_{j}/R^{\prime}}$ and $e^{i\hat{\varphi}_{j,j+1}/R}$ act within $\mathcal{H}_{\mathrm{phy}}$. The compactification radii of $\hat{\theta}_{j}$ and $\hat{\varphi}_{j,j+1}$ are $R^{\prime}$ and $R$, respectively. In terms of these new variables, the Hamiltonian in Eq. (2) becomes
$H_{\mathrm{mV}}=\sum_{j=1}^{N}\left[(\hat{\varphi}_{j+1,j+2}-\hat{\varphi}_{j,j+1})^{2}+(\hat{\theta}_{j+1}-\hat{\theta}_{j})^{2}\right]\,.$ (3)
The continuum limit is taken by sending $N\rightarrow\infty$ and $a\rightarrow 0$ while $L=Na$ is fixed as a constant. After the mapping $\hat{\varphi}_{j+1,j+2}-\hat{\varphi}_{j,j+1}\rightarrow a\partial_{x}\hat{\varphi}(x)$ and $\hat{\theta}_{j+1}-\hat{\theta}_{j}\rightarrow a\partial_{x}\hat{\theta}(x)$, Eq. (3) reduces to the TLL Hamiltonian (1).
Motivated by the geometric structure of crosscap, one should try to construct crosscap states by putting antipodal points into certain maximally entangled states. In the modified Villain model, this line of thought leads to two crosscap states
$|\mathcal{C}_{\pm}\rangle=\mathcal{P}_{G}\prod_{j=1}^{N/2}|\mathcal{B}^{\mathrm{V}}_{\pm}\rangle_{j,j+N/2}\otimes|\mathcal{B}^{\mathrm{L}}_{\mp}\rangle_{j,j+N/2}\,,$ (4)
where
$|\mathcal{B}^{\mathrm{L}}_{\mp}\rangle_{j,j+N/2}$ $=\int_{0}^{2\pi}\mathrm{d}E\,|E\rangle_{j,j+1}|\mp E\rangle_{j+N/2,j+N/2+1}\,,$
|±$|\mathcal{B}^{\mathrm{V}}_{\pm}\rangle_{j,j+N/2}$ $=\int_{-\infty}^{\infty}\mathrm{d}X\,|X\rangle_{j}|\pm X\rangle_{j+N/2}$ (5)
are generalized Bell pairs of antipodal links and sites, respectively. It can be verified that $e^{i(\hat{\theta}_{j}\mp\hat{\theta}_{j+N/2})/R^{\prime}}|\mathcal{C}_{\pm}\rangle=|\mathcal{C}_{\pm}\rangle$ and $e^{i(\hat{\varphi}_{j,j+1}\pm\hat{\varphi}_{j+N/2,j+N/2+1})/R}|\mathcal{C}_{\pm}\rangle=|\mathcal{C}_{\pm}\rangle$ for all $j$ (note that $e^{i\hat{\theta}_{j}/R^{\prime}}$ and $e^{i\hat{\varphi}_{j,j+1}/R}$ commute with $\mathcal{P}_{G}$). In the continuum limit, they become the “sewing conditions”
$e^{i[\hat{\theta}(x)\mp\hat{\theta}(x+L/2)]/R^{\prime}}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle\,,$
$e^{i[\hat{\varphi}(x)\pm\hat{\varphi}(x+L/2)]/R}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle\,,$ (6)
that are imposed on the crosscap states by boundary CFT [52, 53, 54]. It is exhilarating that these conditions can be derived rigorously using our formalism.
For subsequent calculations, the crosscap states may be rewritten using the mode expansion [55]:
$\hat{\theta}(x)$ $=\tilde{\varphi}_{0}+\frac{4\pi x}{L}\pi_{0}+\sum_{k\neq 0}\frac{i}{k}\left[e^{i\frac{2\pi k}{L}x}a_{k}-e^{-i\frac{2\pi k}{L}x}\bar{a}_{k}\right],$
$\hat{\varphi}(x)$ $=\varphi_{0}+\frac{4\pi x}{L}\tilde{\pi}_{0}+\sum_{k\neq 0}\frac{i}{k}\left[e^{i\frac{2\pi k}{L}x}a_{k}+e^{-i\frac{2\pi k}{L}x}\bar{a}_{k}\right]\,,$ (7)
where the zero modes and oscillatory modes satisfy $[\varphi_{0},\pi_{0}]=[\tilde{\varphi}_{0},\tilde{\pi}_{0}]=i$ and $[a_{k},a_{l}]=[\bar{a}_{k},\bar{a}_{l}]=k\delta_{k+l,0}$, respectively. $\varphi_{0}$ ($\tilde{\varphi}_{0}$) is an angular variable with radius $R$ ($R^{\prime}$), and $\pi_{0}$ ($\tilde{\pi}_{0}$) is the associated canonical momentum. The eigenvectors of $\pi_{0}$ and $\tilde{\pi}_{0}$ that can be annihilated by all bosonic annihilation operators $a_{k>0}$ and $\bar{a}_{k>0}$ are denoted as $|n,m\rangle$ ($n,m\in\mathbb{Z}$); they are highest weight states (Virasoro primary states) of the CFT and satisfy $\pi_{0}|n,m\rangle=\tfrac{n}{R}|n,m\rangle$ and $\tilde{\pi}_{0}|n,m\rangle=\tfrac{m}{R^{\prime}}|n,m\rangle$. Using the mode expansion, the compactified boson Hamiltonian (1) becomes
$H=\frac{2\pi v}{L}\left[\pi_{0}^{2}+\tilde{\pi}_{0}^{2}+\sum_{k=1}^{\infty}(a_{-k}a_{k}+\bar{a}_{-k}\bar{a}_{k})-\frac{1}{12}\right],$ (8)
whose energy eigenstates are just Virasoro primary states $|n,m\rangle$ and the descendant states obtained by applying the creation operators $a_{k<0}$ and $\bar{a}_{k<0}$ on top of $|n,m\rangle$. The crosscap states $|\mathcal{C}_{\pm}\rangle$ can be written as
$|\mathcal{C}_{+}\rangle$ $=\kappa_{+}\exp\left[\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]\sum_{n\in\mathbb{Z}}|2n,0\rangle\,,$
$|\mathcal{C}_{-}\rangle$ $=\kappa_{-}\exp\left[-\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]\sum_{m\in\mathbb{Z}}|0,2m\rangle$ (9)
with undetermined overall factors using the mode expansion [56]. The values of $\kappa_{\pm}$ cannot be fixed by normalization. Employing the “cut-and-sew” procedure, we can compare the partition function $Z^{\mathcal{C}}(L,\beta)=\langle\mathcal{C}|e^{-\beta H}|\mathcal{C}\rangle$ on a cylinder with two crosscap boundaries (“tree channel”) and the partition function $Z^{\mathcal{K}}(2\beta,L/2)$ on a Klein bottle (“loop channel”). This analysis yields $\kappa_{+}=\sqrt{R^{\prime}}$ and $\kappa_{-}=\sqrt{R}$ [56].
An appealing property of the crosscap states is that their overlaps with certain highest weight states $|n,m\rangle$ in the CFT are universal numbers depending only on the compactification radius (equivalently, Luttinger parameter):
$|\langle\mathcal{C_{+}}|2n,0\rangle|^{2}=R^{\prime}=\frac{1}{\sqrt{K}},\quad n\in\mathbb{Z}\,,$
$|\langle\mathcal{C_{-}}|0,2m\rangle|^{2}=R=2\sqrt{K},\quad m\in\mathbb{Z}\,.$ (10)
It was found in Ref. [33] that the Luttinger parameter appears in the thermal entropy of compactified boson CFTs on the Klein bottle. In view of the spacetime symmetry of CFTs, it is natural to expect that the same variable can also be found using the ground state alone. Our derivation not only puts this expectation on a firm ground but also provides further insights into crosscap states. Some excited states are incorporated in the present formalism, and perturbations to a pure CFT can be investigated [57].
Spin chains with microscopic U(1) symmetry — To apply the general theory in specific lattice models, the crosscap states should be expressed in terms of the microscopic degrees of freedom. It is helpful to begin with spin-1/2 chains. The first one is the XXZ model
$H_{\mathrm{XXZ}}=\sum_{j=1}^{N}\left(S^{x}_{j}S^{x}_{j+1}+S^{y}_{j}S^{y}_{j+1}+\Delta S^{z}_{j}S^{z}_{j+1}\right)$ (11)
with $\Delta\in(-1,1]$ for which $K=\frac{\pi}{2(\pi-\cos^{-1}\Delta)}$ [1]. The second one is the Haldane-Shastry (HS) model
$H_{\mathrm{HS}}=\left(\frac{\pi}{N}\right)^{2}\sum_{1\leq j<l\leq N}\frac{\mathbf{S}_{j}\cdot\mathbf{S}_{l}}{\sin^{2}\left[\frac{\pi}{N}(j-l)\right]}$ (12)
with $K=1/2$ that provides a perfect lattice realization of the SU(2)${}_{1}$ Wess-Zumino-Witten model [58, 59]. Both models have a U(1) symmetry associated with the conservation of the $z$-component of the total spin.
The success of Eq. (4) suggests that
$|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(|\uparrow\rangle_{j}|\uparrow\rangle_{j+N/2}+|\downarrow\rangle_{j}|\downarrow\rangle_{j+N/2}\right)$ (13)
is likely a crosscap state for spin-1/2 chains with even $N$. However, there is one subtlety in this definition: it corresponds to $|\mathcal{C}_{+}\rangle$ when $\mathrm{mod}(N,4)=0$ but does not give the expected result when $\mathrm{mod}(N,4)=2$. The reason for this can be traced to the constraint $(S_{j}^{z}-S_{j+N/2}^{z})|\mathcal{C}_{\mathrm{latt}}\rangle=0$. Using bosonization, $S^{z}_{j}$ operator reads $S^{z}_{j}\sim\partial_{x}\hat{\theta}(x)+(-1)^{j}\alpha\cos(\hat{\theta}(x)/R^{\prime})$ for XXZ and HS chains, where $\alpha$ is a model-dependent constant. For $\mathrm{mod}(N,4)=0$, the lattice constraint $(S_{j}^{z}-S_{j+N/2}^{z})|\mathcal{C}_{\mathrm{latt}}\rangle=0$ is hence consistent with the constraints of $|\mathcal{C}_{+}\rangle$ in Eq. (6).
We now proceed to demonstrate the validity of Eq. (13). For the XY model (i.e., $\Delta=0$ in the XXZ model) and the HS model, explicit wave functions of many eigenstates are known [60]. This allows us to derive analytical results to support our theory. In both models, the eigenstates that correspond to the Virasoro primary states $|n,0\rangle$ have a Jastrow form in the hardcore boson basis:
$\Psi^{\lambda}(x_{1},\ldots,x_{M})\propto(-1)^{\sum_{l}x_{l}}\prod_{j<k}\left|\sin\frac{\pi(x_{j}-x_{k})}{N}\right|^{\lambda}$ (14)
with $1\leq x_{1}<\cdots<x_{M}\leq N$ denoting the positions of spin-$\uparrow$ sites (with hardcore bosons occupied), and $\lambda=1$ ($\lambda=2$) for the XY (HS) chain. The hardcore boson number $M$ is related to the U(1) charge $n$ via $M=N/2+n$. As $|\mathcal{C}_{\mathrm{latt}}\rangle$ contains only states with even number of bosons, $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M)\rangle$ vanishes identically for odd $M$. This implies $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M=N/2+2n)\rangle=0$ when $\mathrm{mod}(N,4)=2$. As $|\Psi^{\lambda}(M=N/2+2n)\rangle$ corresponds to the Virasoro primary state $|2n,0\rangle$, this indicates $|\mathcal{C}_{\mathrm{latt}}\rangle$ does not realize $|\mathcal{C}_{\pm}\rangle$ in this case [see Eq. (10)]. For $\mathrm{mod}(N,4)=0$, we have $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M)\rangle=0$ for odd $M=N/2+2n+1$ and
$|\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda=1}(M)\rangle|^{2}$ $=1\,,$
$|\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda=2}(M)\rangle|^{2}$ $=\frac{2^{M}(M!)^{3}}{[(M/2)!]^{2}(2M)!}=\sqrt{2}+\mathcal{O}(1/M)\,$ (15)
for even $M=N/2+2n$ [56]. This is in perfect agreement with the field theory prediction in Eq. (10), giving a strong indication that $|\mathcal{C}_{\mathrm{latt}}\rangle$ corresponds to $|\mathcal{C}_{+}\rangle$ for $\mathrm{mod}(N,4)=0$.
Figure 2: Numerically extracted Luttinger parameters in three models. Exact values are indicated using solid or dashed black lines for comparison. (a) The XXZ model with $N=100$ and various $\Delta$. (b) The XXZ model with $\Delta=1$ and various $N$. (c) The next-nearest-neighbor model with $N=100$ and various $J_{2}$. (d) The five-state clock model with various $h$ and $N$.
For the XXZ model with $\Delta{\neq}0$, the ground state can be computed numerically using the density matrix renormalization group (DMRG) [61]. When the ground state $|\Psi\rangle$ is expressed as a matrix product state (MPS) [62, 63], its crosscap overlap $\langle\mathcal{C}_{\rm latt}|\Psi\rangle$ is given by the tensor contraction shown in Fig. 1(b). Numerical results at many different $N$’s and $\Delta$’s are displayed in Fig. 2. The MPS bond dimension is $2000$ for $N=100$ and $6000$ for $N=400$. The accuracy is very good in a wide range of $\Delta$ for a moderate $N=100$. The relative error of $K$ is of the order $10^{-4}$ or smaller when $\Delta\in[-0.95,0.8]$. As $\Delta$ approaches $1$, clear deviation from exact values is observed (not visible in Fig. 2(a) due to its scale) and the relative error reaches $7.20\%$ at $\Delta=1$.
We attribute the obvious deviation at $\Delta=1$ to the marginal terms in the low-energy theory [64], which usually cause strong finite-size effects. The deviation gets smaller as $N$ increases, but the convergence towards the exact value is very slow [see Fig. 2(b)]. To validate our conjecture, we add a next-nearest-neighbor coupling to the XXZ model with $\Delta=1$:
$H_{\rm next}=\sum_{j=1}^{N}(\mathbf{S}_{j}\cdot\mathbf{S}_{j+1}+J_{2}\mathbf{S}_{j}\cdot\mathbf{S}_{j+2})\,.$ (16)
Its low-energy physics contains no marginal term at $J_{2}{\approx}0.24$ [65]. The results for $N=100$ are presented in Fig. 2(c). As $J_{2}$ increases toward $0.24$, the relative error of $K$ decreases and eventually becomes $0.11\%$. This confirms that the marginal terms are responsible for the deviation. Further analysis based on conformal perturbation theory suggests that the Luttinger parameter should be fitted as $K=a[\ln(N)]^{-2}+b$ [57]. The data points with $N\in[100,400]$ in Fig. 2(b) thus yield $K=0.523$ in the thermodynamic limit.
Spin chains without microscopic U(1) symmetry — A fundamental improvement of our method is that microscopic U(1) symmetry is not required. To demonstrate this advantage, we consider the quantum $q$-state clock model [66, 67, 68, 69, 70, 71]
$H_{q{\rm SC}}=-\sum_{j=1}^{N}(\sigma^{{\dagger}}_{j}\sigma_{j+1}+\sigma_{j}\sigma^{{\dagger}}_{j+1})-h\sum_{j=1}^{N}(\tau_{j}+\tau_{j}^{{\dagger}})\,,$ (17)
where the $\mathbb{Z}_{q}$ spin operators are defined by $\sigma|\alpha\rangle=|{\rm mod}(\alpha-1+q,q)\rangle$ and $\tau|\alpha\rangle=e^{i2\pi\alpha/q}|\alpha\rangle$ with $\alpha=1,\ldots,q$. This model is self-dual (SD) at $h=1$. For $q{\geq}5$ and intermediate $h$ (including the SD point), this model hosts a Luttinger liquid that is separated by two Kosterlitz–Thouless (KT) transitions at $h_{\rm KT1}$ and $h_{\rm KT2}$ from gapped phases at small and large $h$ [72, 73]. The transition points are related to each other by duality via $h_{\rm KT1}h_{\rm KT2}=1$. There is no analytical expression for the Luttinger parameter, but its value at the SD point and the KT transition points was predicted to be [68, 34]
$K_{\rm KT1}=\frac{q^{2}}{8},\quad K_{\rm SD}=\frac{q}{2},\quad K_{\rm KT2}=2\,.$ (18)
For the $q$-state clock model, we shall employ the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(\sum_{\alpha=1}^{q}|\alpha\rangle_{j}|\alpha\rangle_{j+N/2}\right)$. Numerical results for the $q=5$ model with $N=50,60,\cdots,90$ and various $h$ are presented in Fig. 2(d). The MPS bond dimension is $1500$ for $N=50$ and $2500$ for $N=90$. At the SD point, finite-size systems are described by CFT with irrelevant perturbations, and the Luttinger parameter has a relative error $0.79\%$ at $N=90$. In the Luttinger liquid regime, the crosscap overlap only exhibits weak dependence on the system size, so the data points for different $N$ almost coincide. For the small and large $h$ phases, the overlap strongly depends on $N$. This picture is supported by Fig. 2(d), but the transition points cannot be located accurately due to the marginal terms at the KT transition points.
Summary and outlook — To summarize, we have demonstrated that the Luttinger parameter can be extracted from the overlaps between individual eigenstates and the crosscap states. The advantages of our method are corroborated by analytical and numerical results. It is also noteworthy that our method does not necessarily require a Hamiltonian, so it may be applied to check if a trial wave function indeed describes the physics of TLL or other CFTs (see, e.g., Refs. [74, 75]). From the theoretical perspective, our results open up an exciting new avenue in extracting conformal data and many interesting directions are yet to be explored. It is clear that the accuracy of our results is strongly affected by marginal terms, so one would like to improve the analytical formula to take their impact into account [57]. The crosscap state used here is constructed from a certain type of Bell states. For integrable spin chains, other types of crosscap states have been studied [47]. Do they have clear counterparts in CFTs? An extension of our method to include multi-component Luttinger liquids, such as the fermionic Hubbard model and SU(3) spin chains, is certainly very desirable. We hope that interesting results on these topics will be reported in the future.
Acknowledgments — We are grateful to Meng Cheng for stimulating discussions. B.Y.T. and Y.H.W. are supported by NNSF of China under Grant No. 12174130. Y.S.Z. is supported by the Sino-German (CSC-DAAD) Postdoc Scholarship Program. W.T. is supported by the Research Foundation Flanders (FWO) via grant GOE1520N. L.W. is supported by the NNSF of China under Grants No. T2225018 and No. 92270107.
Appendix A Appendix A: Tomonaga-Luttinger liquid and compactified boson conformal field theory
The bosonized version of the (single-component) Tomonaga-Luttinger liquid is formulated as a free massless compactified boson field theory in (1+1)-dimensional spacetime. This section is devoted to a brief review of the operator formalism of the compactified boson field theory, and we mostly follow the notation used in Refs. [55, 76]. We assume that the spatial direction is compact with length $L$, and the system admits periodic boundary condition along this direction. Therefore, the spacetime manifold (or “world sheet”), parametrized by coordinates $(x~{}\mathrm{mod}~{}L,t)$, is effectively a cylinder of circumference $L$. The Hamiltonian reads
$H=\frac{v}{8\pi}\int^{L}_{0}\mathrm{d}x\,\left[(\partial_{x}\hat{\varphi})^{2}+(\partial_{x}\hat{\theta})^{2}\right],$ (A1)
where $\hat{\varphi}$ and $\hat{\theta}$ are compactified boson fields, and $v$ is the “speed of light” in this theory. Apart from $v$, the theory is characterized by the compactification radius $R$ of $\hat{\varphi}$, which is related to the Luttinger parameter via $K=R^{2}/4$. The field $\hat{\theta}$ is dual to $\hat{\varphi}$ in the sense that $\partial_{x}\hat{\theta}(x,t)=v^{-1}\partial_{t}\hat{\varphi}(x,t)$.
On this manifold, the mode expansion of the $\hat{\varphi}$-field is given by
$\hat{\varphi}(x,t)=\varphi_{0}+\frac{4\pi}{L}\left(\pi_{0}vt+\tilde{\pi}_{0}x\right)+i\sum_{k\neq 0}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x-vt}{L}}+\bar{a}_{k}e^{-2\pi ik\frac{x+vt}{L}}\right),$ (A2)
where the zero-mode operators satisfy canonical commutation relations (we have formally introduced $\tilde{\varphi}_{0}$ as the operator canonically conjugate to $\tilde{\pi}_{0}$),
$[\varphi_{0},\pi_{0}]=[\tilde{\varphi}_{0},\tilde{\pi}_{0}]=i,$ (A3)
whilst the other modes generate the Heisenberg algebra
$[a_{k},a_{k^{\prime}}]=[\bar{a}_{k},\bar{a}_{k^{\prime}}]=k\delta_{k+k^{\prime}},\quad[a_{k},\bar{a}_{k^{\prime}}]=0.$ (A4)
The fact the $\hat{\varphi}$-field is compactified with radius $R$ means that the field configurations $\hat{\varphi}$ and $\hat{\varphi}+2\pi R$ are identified with each other. This implies that the field can have non-trivial winding around the cylinder, i.e., $\hat{\varphi}(x+L,t)\equiv\hat{\varphi}(x,t)-2\pi mR$ with $m\in\mathbb{Z}$; Eq. (A2) then shows that the operator $\tilde{\pi}_{0}$ has eigenvalues $-mR/2$.
Before proceeding, let us remark on the physical significance of the zero-mode operators $\varphi_{0}$ and $\pi_{0}$. By definition, the “canonical momentum” (density) corresponding to the $\hat{\varphi}$-field is
$\Pi(x,t)\equiv\frac{1}{4\pi v}\partial_{t}\hat{\varphi}(x,t)=\frac{\pi_{0}}{L}+\frac{1}{2L}\sum_{k\neq 0}\left(a_{k}e^{2\pi ik\frac{x-vt}{L}}+\bar{a}_{k}e^{-2\pi ik\frac{x+vt}{L}}\right).$ (A5)
Indeed, the canonical equal-time commutation relation, $[\hat{\varphi}(x,0),\Pi(x^{\prime},0)]=i\delta(x-x^{\prime})$, can readily be verified. The resulting $\mathrm{U}(1)$ conserved charge, $\int_{0}^{L}\mathrm{d}x~{}\Pi(x,t)$, is precisely $\pi_{0}$; this indicates that the latter admits an interpretation as the “centre-of-mass momentum” of the field. In particular, the single-valuedness of the operator $e^{2\pi iR\pi_{0}}$ translating around the compactified dimension (in the space of field configurations; not to be confused with the worldsheet) enforces that the eigenvalues of $\pi_{0}$ are quantized to be $n/R$ with $n\in\mathbb{Z}$. As $\varphi_{0}$ is conjugate to $\pi_{0}$, the former is naturally interpreted as the “centre-of-mass coordinate” operator of the field.
Next, we recombine these zero-mode operators as follows:
$Q$ $=\frac{1}{2}(\varphi_{0}-\tilde{\varphi}_{0}),\quad P=\pi_{0}-\tilde{\pi}_{0};$
$\bar{Q}$ $=\frac{1}{2}(\varphi_{0}+\tilde{\varphi}_{0}),\quad\bar{P}=\pi_{0}+\tilde{\pi}_{0}.$ (A6)
It is easily verified that these new operators still obey the canonical commutation relations,
$[Q,P]=[\bar{Q},\bar{P}]=i.$ (A7)
In terms of this new set of zero-mode operators as well as the “light-cone” coordinates $z\equiv-(x-vt)/L$ and $\bar{z}\equiv-(x+vt)/L$, one arrives at the chiral decomposition of the $\hat{\varphi}$-field, $\hat{\varphi}(x,t)=\phi(z)+\bar{\phi}(\bar{z})$, where
$\phi(z)$ $=Q+2\pi Pz+i\sum_{k\neq 0}\frac{1}{k}a_{k}e^{-2\pi ikz},$
$\bar{\phi}(\bar{z})$ $=\bar{Q}-2\pi\bar{P}\bar{z}+i\sum_{k\neq 0}\frac{1}{k}\bar{a}_{k}e^{2\pi ik\bar{z}}.$ (A8)
Using the relations $\partial_{x}=-L^{-1}(\partial+\bar{\partial})$ and $\partial_{t}=vL^{-1}(\partial-\bar{\partial})$ with $\partial\equiv\partial_{z}$ and ¯$\bar{\partial}\equiv\partial_{\bar{z}}$, one finds that $\hat{\theta}(x,t)=\bar{\phi}(\bar{z})-\phi(z)$, we obtain
$\hat{\theta}(x,t)=\tilde{\varphi}_{0}+\frac{4\pi}{L}\left(\pi_{0}x+\tilde{\pi}_{0}vt\right)-i\sum_{k\neq 0}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x-vt}{L}}-\bar{a}_{k}e^{-2\pi ik\frac{x+vt}{L}}\right).$ (A9)
As $\hat{\theta}(x+L,t)=\hat{\theta}(x,t)+4\pi\pi_{0}$, the quantization of the eigenvalues of $\pi_{0}$ implies that the $\hat{\theta}$-field is also compactified, for which the radius is $2/R$.
Finally, substituting the mode expansions (A2) and (A9) into the expression (A1) of the Hamiltonian, and making use of (A4) as well as the zeta-function regularization (“$1+2+3+\cdots=-1/12$”), one finds after a straightforward computation that
$H=\frac{2\pi v}{L}\left(L_{0}+\bar{L}_{0}-\frac{1}{12}\right),$ (A10)
where
$L_{0}$ $=\frac{1}{2}a_{0}^{2}+\sum_{k=1}^{\infty}a_{-k}a_{k},$
$\bar{L}_{0}$ $=\frac{1}{2}\bar{a}_{0}^{2}+\sum_{k=1}^{\infty}\bar{a}_{-k}\bar{a}_{k},$ (A11)
with $a_{0}\equiv P,~{}\bar{a}_{0}\equiv\bar{P}$, are the zeroth Virasoro generators. The Virasoro primary states (which are normalized by assumption), $|n,m\rangle$, are labelled by the eigenvalues of $a_{0}$ and $\bar{a}_{0}$ as
$a_{0}|n,m\rangle$ $=\left(\frac{n}{R}+\frac{1}{2}mR\right)|n,m\rangle,$
$\bar{a}_{0}|n,m\rangle$ $=\left(\frac{n}{R}-\frac{1}{2}mR\right)|n,m\rangle.$ (A12)
By definition, they are annihilated by the modes $a_{k}$ and $\bar{a}_{k}$ with $k\in\mathbb{Z}^{+}$; other states in the corresponding conformal tower, which are orthogonal to each other and properly normalized, are constructed by acting modes with negative indices on the primary states:
$|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle=\prod_{k=1}^{\infty}\left[\frac{1}{\sqrt{n_{k}!\bar{n}_{k}!}}\left(\frac{a_{-k}}{\sqrt{k}}\right)^{n_{k}}\left(\frac{\bar{a}_{-k}}{\sqrt{k}}\right)^{\bar{n}_{k}}\right]|n,m\rangle,\quad n_{k},\bar{n}_{k}\in\mathbb{Z}^{+}_{0}\,.$ (A13)
One obviously has $|n,m;\{0\},\{0\}\rangle\equiv|n,m\rangle$. This basis trivializes the computation of partition functions, as $L_{0}$ and $\bar{L}_{0}$ act diagonally:
$L_{0}|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle$ $=\left[\frac{1}{2}\left(\frac{n}{R}+\frac{1}{2}mR\right)^{2}+\sum_{k=1}^{\infty}kn_{k}\right]|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle\,,$
$\bar{L}_{0}|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle$ $=\left[\frac{1}{2}\left(\frac{n}{R}-\frac{1}{2}mR\right)^{2}+\sum_{k=1}^{\infty}k\bar{n}_{k}\right]|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle\,.$ (A14)
Appendix B Appendix B: Determining crosscap states from constraints
The mode expansion form of the crosscap states $|\mathcal{C}_{\pm}\rangle$ is given in Eq. (9) of the main text. Below we derive it in two steps:
i) Substitute the mode expansions of the compactified boson fields (A8) into the constraints
$e^{i[\hat{\theta}(x)\mp\hat{\theta}(x+L/2)]/R^{\prime}}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle\,,$
$e^{i[\hat{\varphi}(x)\pm\hat{\varphi}(x+L/2)]/R}|\mathcal{C}_{\pm}\rangle$ $=|\mathcal{C}_{\pm}\rangle$ (A15)
to determine the crosscap states $|\mathcal{C}_{\pm}\rangle$ up to overall factors $\kappa_{\pm}$.
ii) Compare the cylinder partition function with two crosscap boundaries (“tree-channel”), $Z^{\mathcal{C}_{\pm}}(L,\beta)=\langle\mathcal{C}_{\pm}|e^{-\beta H}|\mathcal{C}_{\pm}\rangle$, with the Klein bottle partition function (“loop channel”), $Z^{\mathcal{K}_{\pm}}(2\beta,L/2)=\mathrm{Tr}(\Omega_{\pm}e^{-\beta H})$ to fix the factors $\kappa_{\pm}$.
Step i):
Consider the mode expansions of the compactified boson fields in Eqs. (A2) and (A9) at $t=0$:
$\hat{\theta}(x)=\tilde{\varphi}_{0}+\frac{4\pi x}{L}\pi_{0}-i\sum_{k\neq 0}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x}{L}}-\bar{a}_{k}e^{-2\pi ik\frac{x}{L}}\right),$
$\hat{\varphi}(x)=\varphi_{0}+\frac{4\pi x}{L}\tilde{\pi}_{0}+i\sum_{k\neq 0}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x}{L}}+\bar{a}_{k}e^{-2\pi ik\frac{x}{L}}\right).$ (A16)
To analyze the constraints in Eq. (A15), we consider
$\hat{\theta}(x)-\hat{\theta}(x+L/2)$ $=-2\pi\pi_{0}-2i\sum_{k\neq 0,\,k\;\mathrm{odd}}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x}{L}}-\bar{a}_{k}e^{-2\pi ik\frac{x}{L}}\right),$
$\hat{\varphi}(x)+\hat{\varphi}(x+L/2)$ $=2\varphi_{0}+4\pi\tilde{\pi}_{0}\left(\frac{2x}{L}+\frac{1}{2}\right)+2i\sum_{k\neq 0,\,k\;\mathrm{even}}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x}{L}}+\bar{a}_{k}e^{-2\pi ik\frac{x}{L}}\right),$ (A17)
and
$\hat{\theta}(x)+\hat{\theta}(x+L/2)$ $=2\tilde{\varphi}_{0}+4\pi\pi_{0}\left(\frac{2x}{L}+\frac{1}{2}\right)-2i\sum_{k\neq 0,\,k\;\mathrm{even}}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x}{L}}-\bar{a}_{k}e^{-2\pi ik\frac{x}{L}}\right),$
$\hat{\varphi}(x)-\hat{\varphi}(x+L/2)$ $=-2\pi\tilde{\pi}_{0}+2i\sum_{k\neq 0,\,k\;\mathrm{odd}}\frac{1}{k}\left(a_{k}e^{2\pi ik\frac{x}{L}}+\bar{a}_{k}e^{-2\pi ik\frac{x}{L}}\right).$ (A18)
By inserting these into Eq. (A15) and noting that the constraints (A15) are valid for all $x\in[0,L]$, it is then obvious that the oscillatory part should satisfy
$\left[a_{k}\mp(-1)^{k}\bar{a}_{-k}\right]|\mathcal{C}_{\pm}\rangle$ $=0,\quad k\neq 0\,,$ (A19)
which fixes the form of $|\mathcal{C}_{\pm}\rangle$ to be
$|\mathcal{C}_{\pm}\rangle$ $\propto\exp\left[\pm\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]|\mathcal{C}^{0}_{\pm}\rangle\,.$ (A20)
Here $|\mathcal{C}^{0}_{\pm}\rangle=\sum_{n,m\in\mathbb{Z}}c^{\pm}_{n,m}|n,m\rangle$ encode the zero mode information of the crosscap states, where $c^{\pm}_{n,m}$ are the superposition coefficients that should be determined by using the zero mode constraints from Eq. (A15):
$e^{-2\pi i\pi_{0}/R^{\prime}}|\mathcal{C}^{0}_{+}\rangle=|\mathcal{C}^{0}_{+}\rangle\,,$
$e^{2i\varphi_{0}/R}e^{4\pi i\left(\frac{2x}{L}+\frac{1}{2}\right)\tilde{\pi}_{0}/R}|\mathcal{C}^{0}_{+}\rangle=|\mathcal{C}^{0}_{+}\rangle\,,$ (A21)
and
$e^{2i\tilde{\varphi}_{0}/R^{\prime}}e^{4\pi i\left(\frac{2x}{L}+\frac{1}{2}\right)\pi_{0}/R^{\prime}}|\mathcal{C}^{0}_{-}\rangle=|\mathcal{C}^{0}_{-}\rangle\,,$
$e^{-2\pi i\tilde{\pi}_{0}/R}|\mathcal{C}^{0}_{-}\rangle=|\mathcal{C}^{0}_{-}\rangle\,.$ (A22)
For $|\mathcal{C}^{0}_{+}\rangle=\sum_{n,m\in\mathbb{Z}}c^{+}_{n,m}|n,m\rangle$, the first equation in Eq. (A21) requires $c^{+}_{2n+1,m}=0$, since $e^{-2\pi i\pi_{0}/R^{\prime}}|n,m\rangle=(-1)^{n}|n,m\rangle$ (note that $\pi_{0}|n,m\rangle=\tfrac{n}{R}|n,m\rangle$ and $R^{\prime}=2/R$); the second equation in Eq. (A21) requires $c^{+}_{n,m\neq 0}=0$ and $c^{+}_{2n,0}=c^{+}_{0,0}\;\forall n\in\mathbb{Z}$, since $e^{4\pi i(\frac{2x}{L}+\frac{1}{2})\tilde{\pi}_{0}/R}|n,m\rangle=e^{2\pi i(\frac{2x}{L}+\frac{1}{2})m}|n,m\rangle$ and $e^{2i\varphi_{0}/R}|n,m\rangle=|n+2,m\rangle$ (note that $\tilde{\pi}_{0}|n,m\rangle=-\tfrac{1}{2}mR|n,m\rangle$ and $[\varphi_{0},\pi_{0}]=i$). Thus, we obtain
$|\mathcal{C}^{0}_{+}\rangle\propto\sum_{n\in\mathbb{Z}}|2n,0\rangle\,.$ (A23)
The determination of $|\mathcal{C}^{0}_{-}\rangle$ is similar and will not be repeated. It is given by
$|\mathcal{C}^{0}_{-}\rangle\propto\sum_{m\in\mathbb{Z}}|0,2m\rangle\,.$ (A24)
Altogether, the constraints in Eq. (A15) determine the mode expansion form of the crosscap states
$|\mathcal{C}_{+}\rangle$ $=\kappa_{+}\exp\left[\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]\sum_{n\in\mathbb{Z}}|2n,0\rangle\,,$
$|\mathcal{C}_{-}\rangle$ $=\kappa_{-}\exp\left[-\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}a_{-k}\bar{a}_{-k}\right]\sum_{m\in\mathbb{Z}}|0,2m\rangle$ (A25)
up to overall factors $\kappa_{\pm}$. Without loss of generality, we choose $\kappa_{\pm}\in\mathbb{R}$ and $\kappa_{\pm}>0$ below.
Step ii):
To fix the factors $\kappa_{\pm}$ in Eq. (A25), we first calculate the (“tree channel”) cylinder partition function with crosscap boundary states
$Z^{\mathcal{C}_{\pm}}(L,\beta)=\langle\mathcal{C}_{\pm}|e^{-\beta H}|\mathcal{C}_{\pm}\rangle\,,$ (A26)
where the Hamiltonian $H$ is given in Eq. (A10). The crosscap states in Eq. (A25) can be expanded in the orthonormal basis states in Eq. (A13) as
$|\mathcal{C}_{+}\rangle$ $=\kappa_{+}\prod_{k=1}^{\infty}\sum_{n_{k}=0}^{\infty}\frac{1}{n_{k}!}\left[\frac{(-1)^{k}a_{-k}\bar{a}_{-k}}{k}\right]^{n_{k}}\sum_{n\in\mathbb{Z}}|2n,0\rangle$
$=\kappa_{+}\sum_{n\in\mathbb{Z}}\sum_{n_{1},n_{2},\ldots=0}^{\infty}(-1)^{\sum\limits_{k=1}^{\infty}kn_{k}}|2n,0;\{n_{k}\},\{n_{k}\}\rangle\,,$ (A27)
and
$|\mathcal{C}_{-}\rangle$ $=\kappa_{-}\prod_{k=1}^{\infty}\sum_{n_{k}=0}^{\infty}\frac{1}{n_{k}!}\left[\frac{(-1)^{k+1}a_{-k}\bar{a}_{-k}}{k}\right]^{n_{k}}\sum_{m\in\mathbb{Z}}|0,2m\rangle$
$=\kappa_{-}\sum_{m\in\mathbb{Z}}\sum_{n_{1},n_{2},\ldots=0}^{\infty}(-1)^{\sum\limits_{k=1}^{\infty}(k+1)n_{k}}|0,2m;\{n_{k}\};\{n_{k}\}\rangle\,.$ (A28)
As $|n,m;\{n_{k}\},\{\bar{n}_{k}\}\rangle$ is the Hamiltonian eigenbasis [Eq. (A14)], $Z^{\mathcal{C}_{\pm}}(L,\beta)$ are easy to calculate
$Z^{\mathcal{C}_{+}}(L,\beta)$ $=\kappa_{+}^{2}\sum_{n\in\mathbb{Z}}\sum_{n_{1},n_{2},\ldots=0}^{\infty}\langle 2n,0;\{n_{k}\},\{n_{k}\}|q^{L_{0}+\bar{L}_{0}-1/12}|2n,0;\{n_{k}\},\{n_{k}\}\rangle$
$=\kappa_{+}^{2}\;q^{-1/12}\sum_{n\in\mathbb{Z}}q^{{R^{\prime}}^{2}n^{2}}\sum_{n_{1},n_{2},\ldots=0}^{\infty}q^{2\sum_{k=1}^{\infty}kn_{k}}$
$=\kappa_{+}^{2}\sum_{n\in\mathbb{Z}}q^{{R^{\prime}}^{2}n^{2}}\cdot q^{-1/12}\prod_{k=1}^{\infty}\frac{1}{1-q^{2k}}$
$=\kappa_{+}^{2}\,\frac{\vartheta_{3}(2{R^{\prime}}^{2}\tau)}{\eta(2\tau)}\,,$ (A29)
and similarly
$Z^{\mathcal{C}_{-}}(L,\beta)=\kappa_{-}^{2}\,\frac{\vartheta_{3}(2{R}^{2}\tau)}{\eta(2\tau)}\,,$ (A30)
where $\tau=iv\beta/L$ is the modular parameter and $q=e^{2\pi i\tau}$. The results are expressed in terms of the Dedekind eta function $\eta(\tau)=q^{1/24}\prod_{k=1}^{\infty}(1-q^{k})$ and one of the standard Jacobi’s theta functions $\vartheta_{3}(\tau)=\sum_{k\in\mathbb{Z}}q^{k^{2}/2}$. Under the modular $S$ transformation ($\tau\rightarrow-1/\tau$), Dedekind eta function and Jacobi’s theta function transform as $\sqrt{-i\tau}\vartheta_{3}(\tau)=\vartheta_{3}(-1/\tau)$ and $\sqrt{-i\tau}\eta(\tau)=\eta(-1/\tau)$, respectively. This allows us to rewrite $Z^{\mathcal{C}_{\pm}}$ as
$Z^{\mathcal{C}_{+}}(L,\beta)$ $=\kappa_{+}^{2}\,\frac{1}{R^{\prime}}\,\frac{\vartheta_{3}(-1/2{R^{\prime}}^{2}\tau)}{\eta(-1/2\tau)}\,,$
$Z^{\mathcal{C}_{-}}(L,\beta)$ $=\kappa_{-}^{2}\,\frac{1}{R}\,\frac{\vartheta_{3}(-1/2{R^{\prime}}^{2}\tau)}{\eta(-1/2\tau)}\,.$ (A31)
Now we turn to the (“loop channel”) Klein bottle partition function [33]
$Z^{\mathcal{K}_{\pm}}(2\beta,L/2)=\mathrm{Tr}\left(\Omega_{\pm}e^{-\tilde{H}L/2}\right),$ (A32)
where $\tilde{H}=\frac{2\pi}{\tilde{L}}(L_{0}+\bar{L}_{0}-\frac{1}{12})$ is the compactified boson Hamiltonian [c.f. Eqs. (A1) and (A10)] defined on a circle (in the imaginary time direction) with circumference $\tilde{L}=2v\beta$, which generates evolution along the spatial direction (it can be viewed as the generator of the quantum transfer matrix). The action of the “reflection operator” $\Omega_{\pm}$ on oscillatory modes is given by $\Omega^{-1}_{\pm}a_{k}\Omega_{\pm}=\mp\bar{a}_{k}$ and $\Omega^{-1}_{\pm}\bar{a}_{k}\Omega_{\pm}=\mp a_{k}$ for $k\neq 0$. Applying them to Virasoro primary states gives
$\Omega_{+}|n,m\rangle$ $=|-n,m\rangle\,,$
$\Omega_{-}|n,m\rangle$ $=|n,-m\rangle\,.$ (A33)
Evaluating the Klein bottle partition function $Z^{\mathcal{K}_{\pm}}$ in the Hamiltonian eigenbasis [Eq. (A14)] and noticing that only “left-right-symmetric” states $|0,m;\{n_{k}\},\{n_{k}\}\rangle$ ($|n,0;\{n_{k}\},\{n_{k}\}\rangle$) contribute to $Z^{\mathcal{K}_{+}}$ ($Z^{\mathcal{K}_{-}}$), we obtain
$Z^{\mathcal{K}_{+}}(2\beta,L/2)$ $=\sum_{m\in\mathbb{Z}}\sum_{n_{1},n_{2},\ldots=0}^{\infty}\langle 0,m;\{n_{k}\},\{n_{k}\}|\tilde{q}^{2L_{0}-1/12}|0,m;\{n_{k}\},\{n_{k}\}\rangle$
$=\tilde{q}^{-1/12}\sum_{m\in\mathbb{Z}}\tilde{q}^{m^{2}/{R^{\prime}}^{2}}\sum_{n_{1},n_{2},\ldots=0}^{\infty}\tilde{q}^{2\sum_{k=1}^{\infty}kn_{k}}$
$=\frac{\vartheta_{3}(-1/2{R^{\prime}}^{2}\tau)}{\eta(-1/2\tau)}\,,$ (A34)
and similarly
$Z^{\mathcal{K}_{-}}(2\beta,L/2)$ $=\sum_{n\in\mathbb{Z}}\sum_{n_{1},n_{2},\ldots=0}^{\infty}\langle n,0;\{n_{k}\},\{n_{k}\}|\tilde{q}^{2L_{0}-1/12}|n,0;\{n_{k}\},\{n_{k}\}\rangle$
$=\frac{\vartheta_{3}(-1/2{R}^{2}\tau)}{\eta(-1/2\tau)}$ (A35)
with $\tilde{q}=e^{-\pi L/2v\beta}=e^{-\pi i/2\tau}$.
Considering the spacetime symmetry, the CFT partition function evaluated in tree and loop channels should be equal (“loop channel-tree channel equivalence”), so we expect
$Z^{\mathcal{C}_{\pm}}(L,\beta)=Z^{\mathcal{K}_{\pm}}(2\beta,L/2)\,.$ (A36)
Comparing Eqs. (A31), (A34) and (A35), the factors in the crosscap states (A25) are fixed to be
$\kappa_{+}$ $=\sqrt{R^{\prime}}\,,$
$\kappa_{-}$ $=\sqrt{R}\,.$ (A37)
Appendix C Appendix C: Crosscap state overlap of Jastrow wave functions
In this section, we calculate the overlap between the lattice crosscap state
$|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(|\uparrow\rangle_{j}|\uparrow\rangle_{j+N/2}+|\downarrow\rangle_{j}|\downarrow\rangle_{j+N/2}\right)$ (A38)
and the Jastrow wave function
$|\Psi^{\lambda}(M)\rangle=\sum_{1\leq x_{1}<\cdots<x_{M}\leq N}\Psi^{\lambda}(x_{1},\ldots,x_{M})\sigma_{x_{1}}^{+}\cdots\sigma_{x_{M}}^{+}|\downarrow\cdots\downarrow\rangle$ (A39)
with
$\Psi^{\lambda}(x_{1},\ldots,x_{M})=\mathcal{N}_{N,M,\lambda}\,(-1)^{\sum_{l=1}^{M}x_{l}}\prod_{1\leq j<k\leq M}\left|\sin\frac{\pi(x_{j}-x_{k})}{N}\right|^{\lambda}$ (A40)
for $\mathrm{mod}(N,4)=0$. The number of hardcore bosons, $M$, is taken to be even. For $\lambda=1$ ($\lambda=2$), $|\Psi^{\lambda}(M)\rangle$ is an exact eigenstate of the spin-1/2 XY (Haldane-Shastry) chain. $\mathcal{N}_{N,M,\lambda}$ ensures normalization of $|\Psi^{\lambda}(M)\rangle$ and is chosen to be positive.
To calculate the overlap $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M)\rangle$, we first introduce a unitary $U=\exp\left[i\tfrac{\pi}{2}\sum_{j:\mathrm{odd}}(\sigma^{z}_{j}+1)\right]$ to remove the Marshall sign $(-1)^{\sum_{l=1}^{M}x_{l}}$ in $|\Psi^{\lambda}(M)\rangle$:
$|\tilde{\Psi}^{\lambda}(M)\rangle=U|\Psi^{\lambda}(M)\rangle$ (A41)
with
$\tilde{\Psi}^{\lambda}(x_{1},\ldots,x_{M})$ $=\mathcal{N}_{N,M,\lambda}\,\prod_{1\leq j<k\leq M}\left|\sin\frac{\pi(x_{j}-x_{k})}{N}\right|^{\lambda}\,$
$=\mathcal{N}_{N,M,\lambda}\cdot 2^{-\lambda M(M-1)/2}\prod_{1\leq j<k\leq M}\left|z_{x_{j}}-z_{x_{k}}\right|^{\lambda}\,$
$=\tilde{\mathcal{N}}_{N,M,\lambda}\,\prod_{1\leq j<k\leq M}\left|z_{x_{j}}-z_{x_{k}}\right|^{\lambda}\,,$ (A42)
where $z_{x_{j}}=\exp(i\frac{2\pi}{N}x_{j})$ and $\tilde{\mathcal{N}}_{N,M,\lambda}=2^{-\lambda M(M-1)/2}\mathcal{N}_{N,M,\lambda}$.
For $\mathrm{mod}(N,4)=0$, it is obvious that the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}\rangle$ in Eq. (A38) is invariant under the action of $U$, $U|\mathcal{C}_{\mathrm{latt}}\rangle=|\mathcal{C}_{\mathrm{latt}}\rangle$. The crosscap state overlap is then rewritten as
$\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M)\rangle=\langle\mathcal{C}_{\mathrm{latt}}|U^{{\dagger}}U|\Psi^{\lambda}(M)\rangle=\langle\mathcal{C}_{\mathrm{latt}}|\tilde{\Psi}^{\lambda}(M)\rangle\,.$ (A43)
For even $M$, the overlap $\langle\mathcal{C}_{\mathrm{latt}}|\tilde{\Psi}^{\lambda}(M)\rangle$ can be simplified as follows:
$\langle\mathcal{C}_{\mathrm{latt}}|\tilde{\Psi}^{\lambda}(M)\rangle$ $=\tilde{\mathcal{N}}_{N,M,\lambda}\sum_{1\leq x_{1}<\cdots<x_{M/2}\leq N/2}\Psi^{\lambda}(x_{1},\ldots,x_{M/2},x_{1}+N/2,\ldots,x_{M/2}+N/2)$
$=\tilde{\mathcal{N}}_{N,M,\lambda}\sum_{1\leq x_{1}<\cdots<x_{M/2}\leq N/2}\,\prod_{1\leq j<k\leq M/2}\left|(z_{x_{j}}-z_{x_{k}})(z_{x_{j+N/2}}-z_{x_{k+N/2}})\right|^{\lambda}\cdot\prod_{1\leq j,k\leq M/2}\left|z_{x_{j}}-z_{x_{k+N/2}}\right|^{\lambda}$
$=\tilde{\mathcal{N}}_{N,M,\lambda}\sum_{1\leq x_{1}<\cdots<x_{M/2}\leq N/2}\,\prod_{1\leq j<k\leq M/2}\left|z_{x_{j}}-z_{x_{k}}\right|^{2\lambda}\cdot\prod_{1\leq j<k\leq M/2}\left|z_{x_{j}}+z_{x_{k}}\right|^{2\lambda}\cdot 2^{\lambda M/2}$
$=2^{\lambda M/2}\tilde{\mathcal{N}}_{N,M,\lambda}\sum_{1\leq x_{1}<\cdots<x_{M/2}\leq N/2}\,\prod_{1\leq j<k\leq M/2}\left|z^{2}_{x_{j}}-z^{2}_{x_{k}}\right|^{2\lambda}$
$=2^{\lambda M/2}\tilde{\mathcal{N}}_{N,M,\lambda}\cdot\frac{1}{|\tilde{\mathcal{N}}_{N/2,M/2,\lambda}|^{2}}\,.$ (A44)
For the last step, we have used $\sum_{1\leq x_{1}<\ldots<x_{M/2}\leq N/2}\,\prod_{1\leq j<k\leq M/2}|z_{x_{j}}^{2}-z_{x_{k}}^{2}|^{2}=|\tilde{\mathcal{N}}_{N/2,M/2,\lambda}|^{-2}$, which is the normalization condition of $|\tilde{\Psi}^{\lambda}(M/2)\rangle$ defined on a chain with length $N/2$. Remarkably, the crosscap overlap for the Jastrow wave function $|\Psi^{\lambda}(M)\rangle$ defined in Eqs. (A39) and (A40) is almost solely determined by the normalization factor. Although this interesting result applies to generic $\lambda$, analytical solutions for the normalization factor $\tilde{\mathcal{N}}_{N,M,\lambda}$ are available only for $\lambda=1$ and $2$ (see, e.g., Ref. [60]). For completeness, we provide a self-contained derivation of $\tilde{\mathcal{N}}_{N,M,\lambda}$ for $\lambda=1$ and $2$ below.
C.1 C1: XY chain
For the XY chain with $\lambda=1$, we use the normalization condition $\langle\tilde{\Psi}^{\lambda=1}(M)|\tilde{\Psi}^{\lambda=1}(M)\rangle=1$ to obtain
$|\tilde{\mathcal{N}}_{N,M,\lambda=1}|^{-2}$ $=\sum_{1\leq x_{1}<\cdots<x_{M}\leq N}|\Psi^{\lambda=1}(x_{1},\ldots,x_{M})|^{2}$
$=\sum_{1\leq x_{1}<\cdots<x_{M}\leq N}\prod_{1\leq j<k\leq M}|z_{x_{j}}-z_{x_{k}}|^{2}$
$=\frac{1}{M!}\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\prod_{1\leq j<k\leq M}|z_{x_{j}}-z_{x_{k}}|^{2}\,.$ (A45)
The Jastrow factor $\prod_{1\leq i<j\leq M}(z_{x_{j}}-z_{x_{i}})$ can be written as a Vandermonde determinant
$\prod_{1\leq j<k\leq M}(z_{x_{k}}-z_{x_{j}})=\mathrm{det}\begin{pmatrix}1&z_{x_{1}}&\cdots&z_{x_{1}}^{M-1}\\ 1&z_{x_{2}}&\cdots&z_{x_{2}}^{M-1}\\ \vdots&\vdots&\ddots&\vdots\\ 1&z_{x_{M}}&\cdots&z_{x_{M}}^{M-1}\end{pmatrix}\equiv\mathrm{det}(\mathbb{V})\,,$ (A50)
so we obtain
$|\tilde{\mathcal{N}}_{N,M,\lambda=1}|^{-2}$ $=\frac{1}{M!}\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\mathrm{det}(\mathbb{V})\;\mathrm{det}(\mathbb{V^{*}})$
$=\frac{1}{M!}\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\sum_{\sigma\in S_{M}}\mathrm{sgn}(\sigma)z_{x_{1}}^{\sigma(1)-1}\cdots z_{x_{M}}^{\sigma(M)-1}\sum_{\tau\in S_{M}}\mathrm{sgn}(\tau){(z_{x_{1}}^{*})}^{\tau(1)-1}\cdots{(z_{x_{M}}^{*})}^{\tau(M)-1}$
$=\frac{1}{M!}\sum_{\sigma\in S_{M}}\sum_{\tau\in S_{M}}\mathrm{sgn}(\sigma)\,\mathrm{sgn}(\tau)\left(\sum_{x_{1}=1}^{N}z_{x_{1}}^{\sigma(1)-\tau(1)}\right)\cdots\left(\sum_{x_{M}=1}^{N}z_{x_{M}}^{\sigma(M)-\tau(M)}\right)$
$=\frac{1}{M!}\sum_{\sigma\in S_{M}}\sum_{\tau\in S_{M}}\mathrm{sgn}(\sigma)\,\mathrm{sgn}(\tau)\,\left(N\delta_{\sigma(1),\tau(1)}\right)\cdots\left(N\delta_{\sigma(M),\tau(M)}\right)$
$=\frac{1}{M!}\sum_{\sigma\in S_{M}}N^{M}$
$=N^{M}\,,$ (A51)
where we have used the definition of the determinant [$S_{M}$ is the permutation of $\{1,2,\ldots,M\}$ and $\mathrm{sgn}(\sigma)$ denotes the signature of the permutation $\sigma$] and the identity $\sum_{x=1}^{N}z_{x}^{k}=N\delta_{k,0}$ for $k\in\mathbb{Z}$. Thus, we obtain the normalization factor $\tilde{\mathcal{N}}_{N,M,\lambda=1}=N^{-M/2}$. By using this result and Eq. (A44), we arrive at
$\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda=1}(M)\rangle=2^{M/2}\cdot N^{-M/2}\cdot\frac{1}{[(N/2)^{-M/4}]^{2}}=1\,.$ (A52)
C.2 C2: Haldane-Shastry chain
We now turn to the Haldane-Shastry chain ($\lambda=2$) and determine the normalization factor $\tilde{\mathcal{N}}_{N,M,\lambda=2}$. Using the normalization condition $\langle\tilde{\Psi}^{\lambda=2}(M)|\tilde{\Psi}^{\lambda=2}(M)\rangle=1$, we obtain
$|\tilde{\mathcal{N}}_{N,M,\lambda=2}|^{-2}$ $=\sum_{1\leq x_{1}<\cdots<x_{M}\leq N}|\Psi^{\lambda=2}(x_{1},\ldots,x_{M})|^{2}$
$=\frac{1}{M!}\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\prod_{1\leq j<k\leq M}|z_{x_{j}}-z_{x_{k}}|^{4}$
$=\frac{1}{M!}\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\prod_{1\leq j<k\leq M}(z_{x_{j}}-z_{x_{k}})^{4}\frac{1}{z_{x_{j}}^{2}z_{x_{k}}^{2}}$
$=\frac{1}{M!}\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\prod_{i=1}^{M}z_{x_{i}}^{-2(M-1)}\prod_{1\leq j<k\leq M}(z_{x_{j}}-z_{x_{k}})^{4}\,.$ (A53)
The product $\prod_{1\leq j<k\leq M}(z_{x_{j}}-z_{x_{k}})^{4}$ also has a determinant representation called “confluent alternant” [60]:
$\prod_{1\leq j<k\leq M}(z_{x_{j}}-z_{x_{k}})^{4}=\mathrm{det}\begin{pmatrix}1&z_{x_{1}}&z_{x_{1}}^{2}&z_{x_{1}}^{3}&\cdots&z_{x_{1}}^{2M-1}\\ 0&1&2z_{x_{1}}&3z_{x_{1}}^{2}&\cdots&(2M-1)z_{x_{1}}^{2M-2}\\ 1&z_{x_{2}}&z_{x_{1}}^{2}&z_{x_{1}}^{3}&\cdots&z_{x_{2}}^{2M-1}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&z_{x_{M}}&z_{x_{M}}^{2}&z_{x_{M}}^{3}&\cdots&z_{x_{M}}^{2M-1}\\ 0&1&2z_{x_{M}}&3z_{x_{M}}^{2}&\cdots&(2M-1)z_{x_{M}}^{2M-2}\end{pmatrix}\equiv\mathrm{det}(\mathbb{M})\,.$ (A60)
After expanding the “confluent alternant” $\mathrm{det}(\mathbb{M})$ as a summation of permutation terms (definition of the determinant), the last term in Eq. (A53) can be calculated as follows:
$\phantom{=}\;\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\prod_{i=1}^{M}z_{x_{i}}^{-2(M-1)}\prod_{1\leq j<k\leq M}(z_{x_{j}}-z_{x_{k}})^{4}$
$=\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\prod_{i=1}^{M}z_{x_{i}}^{-2(M-1)}\;\mathrm{det}(\mathbb{M})$
$=\sum_{1\leq x_{1},\ldots,x_{M}\leq N}\prod_{i=1}^{M}z_{x_{i}}^{-2(M-1)}\sum_{\sigma\in S_{2M}}\mathrm{sgn}(\sigma)\,z_{x_{1}}^{\sigma(1)-1}(\sigma(2)-1)z_{x_{1}}^{\sigma(2)-2}\cdots z_{x_{M}}^{\sigma(2M-1)-1}(\sigma(2M)-1)z_{x_{M}}^{\sigma(2M)-2}$
$=\sum_{\sigma\in S_{2M}}\mathrm{sgn}(\sigma)\prod_{j=1}^{M}(\sigma(2j)-1)\prod_{j=1}^{M}\left(\sum_{x_{j}=1}^{N}z_{x_{j}}^{\sigma(2j-1)+\sigma(2j)-(2M+1)}\right)$
$=N^{M}\sum_{\sigma\in S_{2M}}\mathrm{sgn}(\sigma)\prod_{j=1}^{M}(\sigma(2j)-1)\prod_{j=1}^{M}\delta_{\sigma(2j-1)+\sigma(2j),2M+1}\,.$ (A61)
To evaluate the sum, one observes that nonvanishing contributions come from the permutation elements satisfying $\sigma(2j-1)+\sigma(2j)=2M+1$. All such permutation elements are generated from $\sigma=\{1,2M,2,2M-1,\ldots,M,M+1\}$ [which has $\mathrm{sgn}(\sigma)=+1$] by using two types of permutations: i) $\sigma(2j-1)\leftrightarrow\sigma(2j)$ [which has $\mathrm{sgn}(\sigma)=-1$] and ii) $\{\sigma(2j-1),\sigma(2j)\}\leftrightarrow\{\sigma(2l-1),\sigma(2l)\}$ [which has $\mathrm{sgn}(\sigma)=+1$]. Making use of the first type of permutation, the sum can be restricted to $\sigma(2j-1)<\sigma(2j)\;\forall j=1,\ldots,M$ and is then reduced to a simple sum over $\{\sigma(1),\sigma(3),\ldots,\sigma(2M-1)\}\in S_{M}$:
$\phantom{=}\;\sum_{\sigma\in S_{2M}}\mathrm{sgn}(\sigma)\prod_{j=1}^{M}(\sigma(2j)-1)\prod_{j=1}^{M}\delta_{\sigma(2j-1)+\sigma(2j),2M+1}$
$=\sum_{\sigma\in S_{2M},\,\sigma(2j-1)<\sigma(2j)}\mathrm{sgn}(\sigma)\prod_{j=1}^{M}[(\sigma(2j)-1)-(\sigma(2j-1)-1)]\prod_{j=1}^{M}\delta_{\sigma(2j-1)+\sigma(2j),2M+1}$
$=\sum_{\sigma\in S_{2M},\,\sigma(2j-1)<\sigma(2j)}\mathrm{sgn}(\sigma)\prod_{j=1}^{M}[\sigma(2j)-\sigma(2j-1)]\prod_{j=1}^{M}\delta_{\sigma(2j-1)+\sigma(2j),2M+1}$
=$=\sum_{\{\sigma(1),\sigma(3),\ldots,\sigma(2M-1)\}\in S_{M}}(+1)\prod_{j=1}^{M}[2M+1-2\,\sigma(2j-1)]$
$=M!\;(2M-1)(2M-3)\cdots 3\cdot 1$
$=M!\;\frac{(2M)!}{(2M)(2M-2)\cdots 2}$
$=\frac{(2M)!}{2^{M}}\,.$ (A62)
Combining the above results, we obtain
$|\tilde{\mathcal{N}}_{N,M,\lambda=2}|^{-2}=\frac{1}{M!}\cdot N^{M}\cdot\frac{(2M)!}{2^{M}}=\left(\frac{N}{2}\right)^{M}\frac{(2M)!}{M!}\,,$ (A63)
and $\tilde{\mathcal{N}}_{N,M,\lambda=2}=\left(\tfrac{2}{N}\right)^{M/2}\sqrt{\tfrac{M!}{(2M)!}}$. By using this result and Eq. (A44), the crosscap state overlap for the Haldane-Shastry chain is written as
$\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda=2}(M)\rangle$ $=2^{M}\cdot\left(\frac{2}{N}\right)^{M/2}\sqrt{\frac{M!}{(2M)!}}\cdot\left(\frac{N}{4}\right)^{M/2}\frac{M!}{(M/2)!}$
$=\frac{2^{M/2}(M!)^{3/2}}{(M/2)!\sqrt{(2M)!}}\,.$ (A64)
Using the Stirling approximation $M!=\sqrt{2\pi M}\left(\tfrac{M}{e}\right)^{M}\left(1+\tfrac{1}{12M}+\mathcal{O}(M^{-2})\right)$ (valid for $M$ large), we obtain
$\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda=2}(M)\rangle=2^{\frac{1}{4}}\left(1-\frac{1}{16M}+\mathcal{O}\left(\frac{1}{M^{2}}\right)\right)=2^{\frac{1}{4}}+\mathcal{O}(M^{-1}).$ (A65)